OpenStudy (anonymous):

Will give medal!! Step 1:3(2x - 4) = 8 + 2x + 6 Step 2:5x - 7 = 14 + 2 Step 3:5x = 2x = 14 + 7 Step 4:3x = 21 Step 5: x = 7 In which step did the student first make an error and what is the correct step? A) Step 2; 6x - 12 = 14 + 2x B) Step 2; 6x - 7 = 2(6 + x + 4) C) Step 3; 5x - 2x = 14 - 7 D) Step 3; 5x + 2x = 14 + 7

OpenStudy (amistre64):

and you assessment of it is?

OpenStudy (anonymous):

C? @amistre64

OpenStudy (amistre64):

write down step 1 .. can you obtain step 2 from it?

OpenStudy (amistre64):

is step 3 a typo? its got 3 equals signs .. but thats not the first error i see

OpenStudy (amistre64):

***2 equals signs

OpenStudy (anonymous):

Oh my my gosh you're right I'm sorry! Step 1: 3(2x - 4) = 8 + 2x + 6 Step 2: 5x - 7 = 14 + 2x Step 3: 5x - 2x = 12 + 7 Step 4: 3x = 21 Step 5: x = 7 @amistre64

OpenStudy (anonymous):

still 2 errors

OpenStudy (amistre64):

3(2x - 4) = 8 + 2x + 6 now, what would YOU do to process it

OpenStudy (amistre64):

whats your step 2 look like?

OpenStudy (anonymous):

14 became 12 for some reason from step 2 trwp 3

OpenStudy (anonymous):

3(2x) 3(-4) = 8 + 2x + 6 6 -12 = 8 + 2x + 6 ? @americansniper12

OpenStudy (amistre64):

6x-12 on the left, and the right adds to 14+2x is this what they have in their 2nd step?

OpenStudy (anonymous):

Yes @amistre64

OpenStudy (amistre64):

this is really just matching what you have to what they have at this point .. they have: 5x - 7 = 14 + 2x you have: 6x - 12 = 14 + 2x and your assessment you beleive these are alike?

OpenStudy (amistre64):

A) Step 2; 6x - 12 = 14 + 2x B) Step 2; 6x - 7 = 2(6 + x + 4) lets assume you see a difference, and we are to correct it ... what is the correct step 2?

OpenStudy (anonymous):

There were not parentheses? @amister64